OpenStudy (anonymous):

prove the following identity sin^4(x)-cox^4(x)+cos2x=0 kladfalkjflk i just cant get zero its so irritating

ganeshie8 (ganeshie8):

hint : a^2 - b^2 = (a+b)(a-b)

ganeshie8 (ganeshie8):

hint2 : cos^2x - sin^2x = cos(2x)

OpenStudy (anonymous):

i did that alkdjfaldkfj but then i had an extra square on top

ganeshie8 (ganeshie8):

hint3 : cos^2x + sin^2x = 1

OpenStudy (anonymous):

where is the cos^2x + sin^2x = 1? o-o

ganeshie8 (ganeshie8):

\[\large \sin^4(x)-\cos^4(x) = (\sin^2x + \cos^2x)(\sin^2x - \cos^2x)\]

OpenStudy (anonymous):

isn't there a square outside everything?

OpenStudy (anonymous):

a^4-b^4 = (a^2-b^2)^2 doesn't it?

OpenStudy (anonymous):

no, (a+b)^2= a^2+2ab+b^2

ganeshie8 (ganeshie8):

maybe define sin^2x = a and cos^2x = b

OpenStudy (anonymous):

okay so i'll get (a^2-b^2) =(a-b)(a+b) then sub back in (sin^2x-cos^2x)(sin^2x+cos^2x)?

ganeshie8 (ganeshie8):

that looks good

OpenStudy (anonymous):

so (sin^2x-cos^2x) = -cos2x (sin^2x+cos^2x)=1

OpenStudy (anonymous):

therefore answer would be -cos2x+cos2x=0 finally good grief thankyu

ganeshie8 (ganeshie8):

:)